# Decide less, communicate more: On the construct validity of end-to-end fact-checking in medicine

### v2

@@ Appendix I Pilot Claims
Post (r/Diabetes): Affording Medication
so im on a family plan with a 3k/6k out of pocket expense, I think it’s hdhp with a hsa that my husband employer contributes a bit too. I know when I had coverage with my job i had a ppo plan. he’s the one that chooses the plans at his job so im not the best when it comes to explaining the details for it.. was orginally taking metformin but it’s horrible and over the past 2 months it’s been making me sick as a dog so I asked my endocrinologist can I go on something else. she recommended ozempic since alot of patients responded well to it, lost weight, and had a good effect on their sugar. plus it’s taken only weekly in which sounds great for someone like me since I’m not the best with keeping up with medications. back in December since we had hit our 6k deductible I had paid nothing when I recieced the medication so I had no clue what the actual price would be but I nearly had a heart attack when I tried picking it up in the store recently…with my plan I’m at 800 bucks for the thing and optum informed me it’s 2300 (1981 with the discount card) for a 90 day supply. that’s ALOT of money…I was going to purchase farxiga today with optum (1500 dollars) but literally don’t have the money to afford to do so..my car needs a new catalytic converter so finacially I had to make the cut to my medication (my cardiologist put me on that to prevent heart failure since I have "resistant hypertension" that’s not responding well medication) i made a joke to my husband and said I may have to divorce him just so I qualify for that government health insurance. hell looking at it now I may be serious! as a diabetic or anyone in America on any type of medication how are ppl able to afford their insulin/pills/machines/ whatever. our household income is around 85k so there’s not much assistance we can get that im aware of
@@
Post (r/narcolepsy): Do people with IH experience hallucinations?
I am so confused! My MSLT showed IH but my doctor gave me a clinical diagnosis of narcolepsy because I experience hypnopompic hallucinations and sleep paralysis. She told me people with IH dont experience those things which is why she switched the diagnosis. Im confused because Ive read articles that say they are symptoms of IH. I know it doesnt really matter because treatment is the same, but I have this thing in me where I just need to know.
@@
Post (r/CysticFibrosis): Pfizer vaccine
My son, non cf, is having his second pfizer vaccine. He is 25 yrs old. For some reason I’m really nervous about it as he has been told not to exercise for 48hrs afterwards due to heart inflammable young people are getting…obvs this is rare…but my son is extremely active & I’m in a tizz. He’s having now as i write this. I’m extremely proud he is having it as alot of youngsters are refusing it atm but the anxiety over it is making me feel sick.
@@
Post (r/rheumatoidarthritis): Mold and RA
I’m having a bit of a weird issue with mold. I’m currently in the process of being diagnosed with RA. I’ve got achy joints, swelling whole nine yards. I transferred job locations earlier this month and was starting to feel better and my hand swelling finally went down. I then signed up for some overtime in my old job location and after about 2 hours my elbows and hands started to ache and swell. Every time it rains at this building water runs through the walls. I’m certain theirs mold in the walls. Google says long-term toxic mold exposure can mimic RA. Had anyone else had an experience with RA symptoms not ending up being RA or having one large trigger to RA symptoms. After going home and sleeping on things my hands started to feel better but not completely.
@@
Post (r/MultipleSclerosis): copaxone vs aubagio?
My gf is about to switch from once a day copaxon injections to aubagio at the advice of her new neurologist. After doing some research before starting the treatment, she is a bit worried about the liver function concerns with the drug. My gf is bipolar, has high anxiety, and is on several meds for her mental health. I just pulled up a site that compared these 2 drugs and was really angry to see that copaxone patients reported it caused depression, anxiety, and other things the doctor never mentioned. So I am cautiously optimistic that the change is in her best interest. Any thoughts or experiences would be greatly appreciated. My research seems to lean towards the new medication, but we are obviously concerned at least about the liver function monitoring. tyvm
@@
Post (r/thyroidcancer): Calcium supplements (Citracal slow release) and total thyroidectomy Hey all, I’m almost 6 years post total thyroidectomy, and since my providers at the time of my TT didn’t really share any of this info/it was hard to track down, I wanted to put it out there for others.
First– you’ll probably want to get on a calcium supplement. That part I was told. How much calcium I wasn’t told, but eventually found out from a pharmacist to go a bit above the recommended for your age/assigned sex at birth. Normally my recommended would be 1000mg, but because of the TT, it’s 1200mg. Second– wait at least 4 hours after taking your thyroid hormone replacement before taking a calcium supplement. Also was told that, also something everyone here probably already knows. Third– our bodies can only absorb around 500mg of calcium in one go. Fourth– if you’ve had a TT, your body will absorb calcium citrate more effectively than calcium carbonate. I learned this literally a month ago from a PCP who doesn’t specialize in thyroid health, and I’d love to know why my endocrinologist never told me. Now, all of the above led me to be interested in the Citracal slow release, as it’s 1200mg, but released slowly so you can take it once a day and still get all of it. My only issue was that I couldn’t find anywhere that said how long it took to fully release. I was worried that if it took too long, it would prevent my levothyroxine from absorbing the next day. I couldn’t find the answer online, but finally called their questions line today and found out it’s 8 hours. Obviously I’m not a medical provider, I just want more of us to have access to this info, especially since a lot of us have worked with medical providers that don’t give us all of the information we need. Also not here to advertise for that brand specifically, I just wanted something convenient enough to take once a day, and figured others might have had the same question!
Table 14: Ten Pilot Claims. Claim in the post is highlighted. Given population annotation is highlighted in blue. Intervention is highlighted in pink. Outcome is highlighted in green.
Appendix J Refinement Claims
@@
Post (r/POTS): Anyone here with low bp take Ivabradine?
Im just wanting to do a bit of research on different meds before my doctors appointment. Last time they told me they cant medicate me because my bp dropped pretty low during the TTT. However, at rest my bp is normal and even standing up it doesnt drop noticeably low unless Im standing still for a longer period of time. So I just wanted to know if any of yall are in a similar situation and have good (or bad) experiences with this drug. I hear midodrine is good for low bp but its more expensive and the side effects sound kind of iffy to me.
@@
Post (r/Dysthymia): Long-lasting apathetic tendencies, anhedonia etc.
I’m just apathetic in general, and am unable to do even the smallest things. Fluoxetine might have had some positive effects, and I’m supposed to be taking it now, but I can’t even be bothered to get a refill. I can’t tell whether or not my asocial tendencies are a personality trait. I currently have no interest in maintaining a relationship with family or friends. I’ve never been diagnosed with dysthymia - only depression - but a lot of the symptoms seem relevant, and my doctor did mention it at one point.
@@
Post (r/Psychosis): Psychosis and antidepressants
Hey everyone! So some crazy stuff happened to me over the last week. I am on abilify for my psychosis and I have been suffering from depression. My doctor decided to prescribe me Wellbutrin 150mg first. Took it for about five days, started having extreme anxiety and dry mouth. I mentioned this to my doctor and he switched me to Lexapro 5mg. Extreme anxiety and dry mouth but something new happened this time -my fucking delusions and hallucinations came back. I had to legit tell myself my thoughts werent based on reality. But holy crap was it difficult. I didnt take any antidepressants today and already feel better. This is crazy, has anyone experienced anything like this? I didnt think anti depressants would bring out my psychosis. Guess I might have to go the natural route for my depression :(
@@
Post (r/lupus): Cytoxan and prednisone
Rheumatologist says cellcept failed to protect my kidneys and now I have developed lupus nephritis.Im so upset. Prednisone messed up my hips so badly that they both need to be replacedI dont want to get back on it but rheumatologist says its to bring the inflammation down in my kidneys. Ive never been on Cytoxan but the side effects sound identical to a lupus flare. How am I supposed to be positive with news like this? I feel so defeatedI dont know what to do.
@@
Post (r/Diabetes): Can metformin replace insulin?
I realize this is definitely case-by-case but Im curious to know if anyone has been able to get off of insulin and take just metformin? When I was diagnosed with type 2 I was automatically put on insulin and generally take quite a bit of it, but now after some research Im considering asking my doctor to try treatment through metformin in February.
Table 15: Retrieval Test Ratings (Population, Intervention, Outcome, Overall) by Claim. Full Score is 24. Claim in the post is highlighted. Given population annotation is highlighted in blue. Intervention is highlighted in pink. Outcome is highlighted in green.
Appendix K Automatic Pipeline for Claim RCT Verifiability
We first started with this prompt for RCT Verifiability.
Can the following claim from the text from r/subreddit be verified by conducting a randomized controlled trial? To be verifiable the claim must have a clear intervention that would be ethical to perform given the context. Give a classification of verifiable or not verifiable.
Claim: [claim]
Text from r/[subreddit]: [post]
On a small set, our medical expert did not approve.
Our medical expert proposed this refined prompt with a persona-style instruction to be a potential clinical trialist as well as more detailed ethical guidelines:
Claim: [claim]
Text from r/[subreddit]: [post]
Appendix L Implementation
Appendix M Plain Language Explanation Guideline
  • (Optional) Medical Addendum at end.
Appendix N Changes between annotation rounds
In the first round, ten claims were annotated without PIO claim contextualization and the expert support field. In the second round, we adjusted the annotation guidelines to clarify labels to increase agreement. The resulting agreement was still low. After receiving feedback from annotators, we revamped our system by adding a filter to filter non-RCT verifiable claims, extract and provide PIO elements for the annotators to focus on, improved the retrieval system, and further clarified the annotation guidelines. For the third round, experts annotated five claims with these new updates.
Appendix O Ethical Concerns
The posts found within the RedHOT dataset do contain health-related comments that are inherently sensitive. To respect this sensitivity, the authors of the RedHOT dataset notified all users of their inclusion in this dataset and provided them with the opportunity to opt-out. They also did not release the data directly, but instead provided a script to download content from Reddit so that individuals may be able to remove their post in the future. In this work, we do directly release a small subset of these posts from RedHOT that we used in our annotation study. In our released data, we do not reveal the username of the author of the post. We only include the text from the post and information about the subreddit in which it was found. Considering the measures taken by the authors of the RedHOT dataset and the fact that these posts have been publicly available on Reddit for at least more than an year, we believe it is safe to publicly release this data.
We have consulted with our Institutional Review Board (IRB) about the nature of our work and confirmed that the use of the RedHOT data and the subsequent annotation study using these data do not constitute research of human subjects. However, we do acknowledge that certain uses of this data may be considered sensitive. We strongly encourage researchers to obtain prior approval from their own IRB regarding the intended use of the data released by this work.
Appendix P License Information
@@
Appendix Q Annotation Interface
We used a web-based annotation interface to collect annotations from our expert medical annotators. Figure 3 shows how we present a claim with its surrounding context, extracted PIO elements for that claim, and retrieved abstracts corresponding to that claim. In this setup, we highlight the claim within the post in which its found along with any PIO spans as determined by data found in the RedHOT dataset. We also present extracted PIO elements (See Appendix E) in a separate box with a rewritten claim created by inputting these elements in a template. All of this information was provided for the benefit of the annotator to clearly understand the claim in question.

### v3

@@ Appendix I PIO Extraction
Post (r/Diabetes): Affording Medication so im on a family plan with a 3k/6k out of pocket expense, I think it’s hdhp with a hsa that my husband employer contributes a bit too. I know when I had coverage with my job i had a ppo plan. he’s the one that chooses the plans at his job so im not the best when it comes to explaining the details for it.. was orginally taking metformin but it’s horrible and over the past 2 months it’s been making me sick as a dog so I asked my endocrinologist can I go on something else. she recommended ozempic since alot of patients responded well to it, lost weight, and had a good effect on their sugar. plus it’s taken only weekly in which sounds great for someone like me since I’m not the best with keeping up with medications. back in December since we had hit our 6k deductible I had paid nothing when I recieced the medication so I had no clue what the actual price would be but I nearly had a heart attack when I tried picking it up in the store recently…with my plan I’m at 800 bucks for the thing and optum informed me it’s 2300 (1981 with the discount card) for a 90 day supply. that’s ALOT of money…I was going to purchase farxiga today with optum (1500 dollars) but literally don’t have the money to afford to do so..my car needs a new catalytic converter so finacially I had to make the cut to my medication (my cardiologist put me on that to prevent heart failure since I have "resistant hypertension" that’s not responding well medication) i made a joke to my husband and said I may have to divorce him just so I qualify for that government health insurance. hell looking at it now I may be serious! as a diabetic or anyone in America on any type of medication how are ppl able to afford their insulin/pills/machines/ whatever. our household income is around 85k so there’s not much assistance we can get that im aware of
@@
Post (r/narcolepsy): Do people with IH experience hallucinations? I am so confused! My MSLT showed IH but my doctor gave me a clinical diagnosis of narcolepsy because I experience hypnopompic hallucinations and sleep paralysis. She told me people with IH dont experience those things which is why she switched the diagnosis. Im confused because Ive read articles that say they are symptoms of IH. I know it doesnt really matter because treatment is the same, but I have this thing in me where I just need to know.
@@
Post (r/CysticFibrosis): Pfizer vaccine My son, non cf, is having his second pfizer vaccine. He is 25 yrs old. For some reason I’m really nervous about it as he has been told not to exercise for 48hrs afterwards due to heart inflammable young people are getting…obvs this is rare…but my son is extremely active & I’m in a tizz. He’s having now as i write this. I’m extremely proud he is having it as alot of youngsters are refusing it atm but the anxiety over it is making me feel sick.
@@
Post (r/rheumatoidarthritis): Mold and RA I’m having a bit of a weird issue with mold. I’m currently in the process of being diagnosed with RA. I’ve got achy joints, swelling whole nine yards. I transferred job locations earlier this month and was starting to feel better and my hand swelling finally went down. I then signed up for some overtime in my old job location and after about 2 hours my elbows and hands started to ache and swell. Every time it rains at this building water runs through the walls. I’m certain theirs mold in the walls. Google says long-term toxic mold exposure can mimic RA. Had anyone else had an experience with RA symptoms not ending up being RA or having one large trigger to RA symptoms. After going home and sleeping on things my hands started to feel better but not completely.
@@
Post (r/MultipleSclerosis): copaxone vs aubagio? My gf is about to switch from once a day copaxon injections to aubagio at the advice of her new neurologist. After doing some research before starting the treatment, she is a bit worried about the liver function concerns with the drug. My gf is bipolar, has high anxiety, and is on several meds for her mental health. I just pulled up a site that compared these 2 drugs and was really angry to see that copaxone patients reported it caused depression, anxiety, and other things the doctor never mentioned. So I am cautiously optimistic that the change is in her best interest. Any thoughts or experiences would be greatly appreciated. My research seems to lean towards the new medication, but we are obviously concerned at least about the liver function monitoring. tyvm
@@
Post (r/thyroidcancer): Calcium supplements (Citracal slow release) and total thyroidectomy Hey all, I’m almost 6 years post total thyroidectomy, and since my providers at the time of my TT didn’t really share any of this info/it was hard to track down, I wanted to put it out there for others. First– you’ll probably want to get on a calcium supplement. That part I was told. How much calcium I wasn’t told, but eventually found out from a pharmacist to go a bit above the recommended for your age/assigned sex at birth. Normally my recommended would be 1000mg, but because of the TT, it’s 1200mg. Second– wait at least 4 hours after taking your thyroid hormone replacement before taking a calcium supplement. Also was told that, also something everyone here probably already knows. Third– our bodies can only absorb around 500mg of calcium in one go. Fourth– if you’ve had a TT, your body will absorb calcium citrate more effectively than calcium carbonate. I learned this literally a month ago from a PCP who doesn’t specialize in thyroid health, and I’d love to know why my endocrinologist never told me. Now, all of the above led me to be interested in the Citracal slow release, as it’s 1200mg, but released slowly so you can take it once a day and still get all of it. My only issue was that I couldn’t find anywhere that said how long it took to fully release. I was worried that if it took too long, it would prevent my levothyroxine from absorbing the next day. I couldn’t find the answer online, but finally called their questions line today and found out it’s 8 hours. Obviously I’m not a medical provider, I just want more of us to have access to this info, especially since a lot of us have worked with medical providers that don’t give us all of the information we need. Also not here to advertise for that brand specifically, I just wanted something convenient enough to take once a day, and figured others might have had the same question!
Table 10: Pilot claims.
Appendix N Refinement Claims
@@
Post (r/POTS): Anyone here with low bp take Ivabradine? Im just wanting to do a bit of research on different meds before my doctors appointment. Last time they told me they cant medicate me because my bp dropped pretty low during the TTT. However, at rest my bp is normal and even standing up it doesnt drop noticeably low unless Im standing still for a longer period of time. So I just wanted to know if any of yall are in a similar situation and have good (or bad) experiences with this drug. I hear midodrine is good for low bp but its more expensive and the side effects sound kind of iffy to me.
@@
Post (r/Dysthymia): Long-lasting apathetic tendencies, anhedonia etc. I’m just apathetic in general, and am unable to do even the smallest things. Fluoxetine might have had some positive effects, and I’m supposed to be taking it now, but I can’t even be bothered to get a refill. I can’t tell whether or not my asocial tendencies are a personality trait. I currently have no interest in maintaining a relationship with family or friends. I’ve never been diagnosed with dysthymia - only depression - but a lot of the symptoms seem relevant, and my doctor did mention it at one point.
@@
Post (r/Psychosis): Psychosis and antidepressants Hey everyone! So some crazy stuff happened to me over the last week. I am on abilify for my psychosis and I have been suffering from depression. My doctor decided to prescribe me Wellbutrin 150mg first. Took it for about five days, started having extreme anxiety and dry mouth. I mentioned this to my doctor and he switched me to Lexapro 5mg. Extreme anxiety and dry mouth but something new happened this time -my fucking delusions and hallucinations came back. I had to legit tell myself my thoughts werent based on reality. But holy crap was it difficult. I didnt take any antidepressants today and already feel better. This is crazy, has anyone experienced anything like this? I didnt think anti depressants would bring out my psychosis. Guess I might have to go the natural route for my depression :(
@@
Post (r/lupus): Cytoxan and prednisone Rheumatologist says cellcept failed to protect my kidneys and now I have developed lupus nephritis.Im so upset. Prednisone messed up my hips so badly that they both need to be replacedI dont want to get back on it but rheumatologist says its to bring the inflammation down in my kidneys. Ive never been on Cytoxan but the side effects sound identical to a lupus flare. How am I supposed to be positive with news like this? I feel so defeatedI dont know what to do.
@@
Post (r/Diabetes): Can metformin replace insulin? I realize this is definitely case-by-case but Im curious to know if anyone has been able to get off of insulin and take just metformin? When I was diagnosed with type 2 I was automatically put on insulin and generally take quite a bit of it, but now after some research Im considering asking my doctor to try treatment through metformin in February.
Table 11:
Appendix O License Information